OpenStudy (anonymous):

What is the measure of the angle between v = <4, 0> and w = <6, 1>? 9.5° 16.3° 18.9° 21.4°

OpenStudy (amorfide):

\[u.v=|u||v|\cos \theta\]

OpenStudy (michele_laino):

hint: we can apply this formula: \[\Large \cos \theta = \frac{{u \cdot v}}{{\left\| u \right\|\;\left\| v \right\|}} = ...?\]

OpenStudy (michele_laino):

where \theta is the requested angle

OpenStudy (anonymous):

so the answer is C

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have: \[\cos \theta = \frac{{u \cdot v}}{{\left\| u \right\|\;\left\| v \right\|}} = \frac{{24}}{{25}}\] so: \[\theta = \arccos \left( {\frac{{24}}{{25}}} \right) = ...?\]

OpenStudy (anonymous):

16.3

OpenStudy (anonymous):

but I wonder how you get the 25 at the bottom?

OpenStudy (michele_laino):

I'm sorry I have made an error, since we can write: \[\begin{gathered} \left\| u \right\| = length\;of\;u = 4 \hfill \\ \left\| v \right\| = length\;of\;v = \sqrt {{6^2} + {1^2}} = \sqrt {36 + 1} = \sqrt {37} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so: \[\begin{gathered} \cos \theta = \frac{{u \cdot v}}{{\left\| u \right\|\;\left\| v \right\|}} = \frac{{24}}{{4\sqrt {37} }} \hfill \\ \hfill \\ \theta = \arccos \left( {\frac{{24}}{{4\sqrt {37} }}} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

sorry again for my error

OpenStudy (anonymous):

you straight

OpenStudy (anonymous):

hey got it wrong but it is A

OpenStudy (michele_laino):

yes! it is A

OpenStudy (michele_laino):

\[\theta = \arccos \left( {\frac{{24}}{{4\sqrt {37} }}} \right) = 9.46\]